OpenStudy (anonymous):

How to simplify: (1-3i)(2+i)/(2-i) ?

OpenStudy (luigi0210):

\[\frac{ (1-3i)(2+i) }{ 2-i }\]

OpenStudy (luigi0210):

Is that it?

OpenStudy (anonymous):

yes! @Luis_Rivera

OpenStudy (luigi0210):

Well then all you do is FOIL the top

OpenStudy (anonymous):

I got 5(I+1)/(2-i) so far

OpenStudy (anonymous):

@Luigi0210 would this be one of those situtations when you may mulitply by the reciprical "2+i"..?

OpenStudy (luigi0210):

You could just leave it as \[\frac{ 5-5i }{ 2-i }\]

OpenStudy (luigi0210):

*conjugate, and yes sir!

OpenStudy (luigi0210):

So just as viper says, multiply the top and bottom by the conjugate of 2+i

OpenStudy (anonymous):

@Luigi0210 Are you sure you cant simplify it because my teacher did (?... (-2) ) .. He took off points for not simplifying

OpenStudy (anonymous):

OH I just saw that okay so 5(I+1)(2+i)

OpenStudy (luigi0210):

\[\frac{ 5-5i }{ 2-i }*\frac{ 2+i }{ 2+i }=\frac{ 10+5i-10i-5i^2 }{ 4-i^2 }\] Can you finish it from here Miss Princess?

OpenStudy (anonymous):

(1 - 3i) x 2 + i / 2 - i = 3 - 1

OpenStudy (anonymous):

@Luigi0210 Oh sorry I forgot! Yea ..

OpenStudy (anonymous):

then youd combine like terms

OpenStudy (anonymous):

would you expand the denom.

OpenStudy (luigi0210):

\[i^2=-1\]

OpenStudy (anonymous):

Did you get -5i+15/5 ??

OpenStudy (luigi0210):

Yup, and simplified down farther it is 3-i.

OpenStudy (luigi0210):

*or -i+3

OpenStudy (anonymous):

Haha yea it's 3 - i

OpenStudy (anonymous):

The works up there sorry I don't want to post it again haha!

OpenStudy (luigi0210):

What up pooper? :D

OpenStudy (anonymous):

Nothing I'm just bored so I decided to help people with math :/ haha Marios bro

OpenStudy (luigi0210):

Ah, interesting I know :P

OpenStudy (anonymous):

Haha!

OpenStudy (anonymous):

I'm going to continue on now! :D

OpenStudy (luigi0210):

ha, well alright.. good day to you :P

OpenStudy (anonymous):

thnx!

OpenStudy (anonymous):

You too luigi! :*

OpenStudy (luigi0210):

No problem :P Enjoy the rest of your day you two :P